Please help me
For the function f(x)=x^2-2x+1
(a) find f(0)
(b) solve f(x)=0

Answer by venugopalramana(3286)

Please help me
For the function f(x)=x^2-2x+1
(a) find f(0)=0^2-2*0+1=1
(b) solve f(x)=0=X^2-X-X+1=X(X-1)+1(X-1)=(X-1)(X-1)
HENCE X-1=0....OR....X=1